### v1

# Machine Learning for Two-Sample Testing under Right-Censored Data: A Simulation Study

Abstract
The focus of this study is to evaluate the effectiveness of Machine Learning (ML) methods for two-sample testing with right-censored observations. To achieve this, we develop several ML-based methods with varying architectures and implement them as two-sample tests. Each method is an ensemble (stacking) that combines predictions from classical two-sample tests. This paper presents the results of training the proposed ML methods, examines their statistical power compared to classical two-sample tests, analyzes the distribution of test statistics for the proposed methods when the null hypothesis is true, and evaluates the significance of the features incorporated into the proposed methods. All results from numerical experiments were obtained from a synthetic dataset generated using the Smirnov transform (Inverse Transform Sampling) and replicated multiple times through Monte Carlo simulation. To test the two-sample problem with right-censored observations, one can use the proposed two-sample methods. All necessary materials (source code, example scripts, dataset, and samples) are available on GitHub and Hugging Face.
keywords:
Hypothesis Testing; Two-Sample Problem; Lifetime Data; Machine Learning; Monte-Carlo Simulation; GitHub.
articletype: ARTICLE
1 Introduction
Two-sample problem testing (and hypothesis testing in general) on censored data is an integral part of solving problems in reliability, physics, biology, medicine, economics, bioinformatics, sociology, engineering, and many other fields [1]. It is well known that there is no universal method preferred in all usage scenarios for this class of problems. Given the wide applicability, there is a constant need for improvement in the methods used. Machine learning methods, which are currently being applied in various domains, can help address this need [2].
In the construction of classical two-sample tests, reliance is primarily placed on certain parametric or non-parametric assumptions [3]. This allows the method to effectively distinguish between certain competing hypotheses, but unfortunately, it may struggle to differentiate others. Researchers have made numerous attempts to combine the properties of different tests [4, 5, 6]. However, such attempts are complicated by the peculiarities of the subject area, which involves censored data. In the literature, there are also modern works that contrapose ML methods against classical two-sample tests [7] or for the uncensored case [8, 9, 10]. However, these comparisons often involve a small number of methods. In our work, we propose considering a wide range of diverse two-sample tests. Each two-sample tests is treated as an feature, and through the use of ML-methods, we create a stacking (ensemble) that combines these features.
Therefore, the main objectives of this research are as follows:
  1. 1.
    Develop methods for two-sample problem under right-censored data using ML-methods;
  2. 2.
    Evaluate the effectiveness of the proposed methods compared to existing classical methods;
  3. 3.
    Describe the practical applicability of the proposed methods.
  4. 4.
    Publish the repositories and results of the proposed methods for public use.
2 Related Works
As mentioned earlier, existing two-sample tests often rely on constructing a specific sensitive statistic based on certain parametric or non-parametric assumptions. This approach enhances sensitivity to certain types of alternatives.
One group of such tests includes various generalizations of classical tests such as the Cramér–von-Mises and Kolmogorov-Smirnov for censored data, as shown in the works of Gill [11], Fleming et al. [12], Koziol [13], Schumacher [14], and Fleming [15].
Another group of tests is based on the use of weight functions, such as the weighted log-rank test. The use of different weight functions generates a multitude of different two-sample tests with new properties compared to the original ones. For example, Gehan [16], Peto and Peto [17], Tarone and Ware [18], Prentice [19], Kalbfleisch and Prentice [20], Fleming and Harrington [21, 22].
The third group of tests is based on the assumption of a specific form of competing hypotheses. This approach was used in the works of Bagdonavičius and Nikulin. They assumed that competing hypotheses could correspond to models with single (SCE-model [23]) or multiple (MCE-model [24]) intersections, and the GPH-model [25].
Additionally, due to their simplicity and efficiency, the method of constructing methods using auxiliary functions has gained popularity. These methods involve computation the value of the selection function in advance, and then using this value to make decisions about the value of the test statistic. For example, Martinez and Naranjo [5], as well as Philonenko and Postovalov [26], propose using the selection function to choose between the log-rank and Wilcoxon tests. In another work [27], authors combine the Bagdonavičius-Nikulin and weighted Kaplan-Meier tests by the selection function.
Established the potential effectiveness of this approach, the question arises of constructing a selection function that is more sensitive to the characteristics of the underlying tests. ML methods can be helpful in this regard, as they can be considered as a selection function on a set of two-sample tests, using each test as a feature in the model.
Thus, in Section 3, we present the problem formulation and describe existing statistical methods. In Section 4, we describe the proposed methods for the two-sample problem using machine learning. In Section 5, we provide the results of numerical experiments, including the training results of the proposed methods, power study of the proposed methods, study of the test statistic distribution, and feature importance results.
3 Materials & Methods
3.1 Problem Statement
Suppose that we have two samples of continues variables $\xi_{1}$ and $\xi_{2}$ respectively, $X_{1}=\{t_{11},t_{12},\cdots,t_{1n_{1}}\}$ and $X_{2}=\{t_{21},t_{22},\cdots,t_{2n_{2}}\}$ of two survival distributions $S_{1}(t)$ and $S_{2}(t)$. The observation $t_{ij}=min(T_{ij},C_{ij})$, where $T_{ij}$ and $C_{ij}$ are the failure and censoring times for the $j$th object of the $i$th group. $T_{ij}$ and $C_{ij}$ are i.i.d. with a cumulative distribution function (CDF) $F_{i}(t)$ and $F_{i}^{C}(t)$ respectively. Survival curve means the probability of survival in the time interval $(0,t)$, i.e. $S_{i}(t)=P\{\xi_{i}>t\}=1-F_{i}(t).$
Then the null hypothesis is $H_{0}:S_{1}(t)=S_{2}(t)$ against an alternative hypothesis $H_{A}:S_{1}(t)\neq S_{2}(t)$.
Further, we suppose that the elements of samples are ordered: $t_{11}<\cdots<t_{1n_{1}}$ and $t_{21}<\cdots<t_{2n_{2}}$. Denote the censoring indicators $c_{ij}$ as $c_{ij}=\begin{cases}0,t_{ij}\text{~{}is failure,}\\ 1,t_{ij}\text{~{}is censored,}\\ \end{cases}$ and the censoring rate $r_{i}$ as $r_{i}=\sum_{j=1}^{n_{i}}c_{ij}/n_{i}$.
3.2 Two-Sample Tests
3.2.1 Log-rank test
The log-rank test is also called as Mantel–Haenszel test or Mantel–Cox test [3, 28, 29]. The test statistic is
$S_{LG}=\frac{\sum_{i=1}^{n}(1-c_{i})-\sum_{j=1}^{i}\frac{1}{n-j+1}}{\sqrt{(\sum_{i=1}^{n}(1-c_{i})\frac{n-i}{n-i+1})\frac{n_{1}n_{2}}{n(n-1)}}}.$
The null hypothesis $H_{0}$ is rejected with the significance level $\alpha$, if $\left|S_{LG}\right|>z_{1-\alpha/2}$, where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution.
3.2.2 Generalizations of Wilcoxon test
Gehan’s generalized Wilcoxon test. This test is also called as Gehan-Breslow-Wilcoxon test. The test statistic [3, 16] is
$S_{G}=\frac{\sum_{i=1}^{n}(1-\nu_{i})h_{i}}{\sqrt{\frac{n_{1}n_{2}}{n(n-1)}\sum_{i-1}^{n}(1-\nu_{i})h_{i}^{2}}},$
where $h_{i}=\sum_{j=1}^{n}\nu_{j}h_{ij}$ and $h_{ij}=\begin{cases}+1,t_{i}>t_{j}~{}\&~{}c_{j}=0~{}\&~{}\nu_{i}=0~{}\&~{}\nu_{j}=1,\\ -1,t_{i}<t_{j}~{}\&~{}c_{j}=0~{}\&~{}\nu_{i}=0~{}\&~{}\nu_{j}=1,\\ 0,\text{~{}otherwise.}\\ \end{cases}$
The null hypothesis $H_{0}$ is rejected with the significance level $\alpha$, if $\left|S_{G}\right|>z_{1-\alpha/2}$, where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution.
Peto and Peto’s generalized Wilcoxon test. The test statistic [3, 17] is
$S_{P}=\frac{n(n-1)\sum_{i=1}^{n}u_{i}(1-\nu_{i})}{n_{1}n_{2}\sum_{i=1}^{n}u_{i}^{2}},$
where $u_{i}=\begin{cases}\hat{S}_{KM}(t_{i})+\hat{S}_{KM}(t_{k_{i}}),c=0,\\ \hat{S}_{KM}(t_{k_{i}})-1,c=1,\\ \end{cases}$ $k_{i}=max\{j|j\in\{0,...,i-1\},c_{j}=0\}$, $c_{0}=0$ and $\hat{S}_{KM}(t)$ is the Kaplan-Meier estimator [30] of $S(t)$ by the pooled sample $T$.
The null hypothesis $H_{0}$ is rejected with the significance level $\alpha$, if $\left|S_{P}\right|>z_{1-\alpha/2}$, where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$-quantile of the standard normal distribution.
3.2.3 Weighted tests
Weighted log-rank test. Based on the findings of [31], we describe the statistics in a unified format, covered all previously presented tests and incorporating additional ones. The test statistic can be computed as $S_{WLG}=U^{2}/\sigma$ where
$U=\sum_{j=1}^{n_{1}}(1-c_{1j})K(t_{1j})\frac{Y_{2}(t_{1j})}{Y(t_{1j})}+\sum_{j=1}^{n_{2}}(1-c_{2j})K(t_{2j})\frac{Y_{1}(t_{2j})}{Y(t_{2j})},$
$\sigma=\sum_{i=1}^{2}\sum_{j=1}^{n_{i}}(1-c_{ij})K^{2}(t_{ij})\frac{Y_{1}(t_{ij})Y_{2}(t_{ij})}{Y^{2}(t_{ij})},$
$Y(t)=Y_{1}(t)+Y_{2}(t),Y_{i}(t)=\sum_{j=1}^{n_{i}}Y_{ij}(t),Y_{ij}=1_{t_{ij}\geq t},i=1,2.$
Using a certain parameterizations of $K(t)$, it is possible to vary a behaviour of the test statistic. We present some possible examples of parameterization:
  1. 1.
    log-rank (Mantel–Haenszel) statistic: $K(t)=1/\sqrt{n_{1}+n_{2}}$;
  2. 2.
    Gehan-Breslow-Wilcoxon statistic: $K(t)=Y(t)/\sqrt{(n_{1}+n_{2})^{3}}$;
  3. 3.
    Peto-Peto-Prentice statistic: $K(t)=\hat{S}(t)/\sqrt{n_{1}+n_{2}}$;
  4. 4.
    Tarone-Ware statistic: $K(t)=\sqrt{Y(t)}/(n_{1}+n_{2})$;
  5. 5.
    Prentice statistic: $K(t)=\frac{\hat{S}(t)}{\sqrt{n_{1}+n_{2}}}\cdot\frac{Y(t)}{Y(t)+1}$;
where $\hat{S}(t)$ is the Kaplan-Meier estimator on the pooled sample.
The test statistic $S_{WLG}$ is asymptotically distributed as chi-square distribution with one degree of freedom. The test statistic has a right-critical area.
Weighted Kaplan-Meier test. The weighted Kaplan–Meier test statistic measures the average difference between the survival functions.
Let $\hat{S}_{1}(t)$ and $\hat{S}_{2}(t)$ are the Kaplan–Meier estimates [30] of the $S_{1}(t)$ and $S_{2}(t)$ consequently, and $\hat{S}(t)$ is the Kaplan–Meier estimate of the pooled sample. For every sample $X_{i}$, we construct additional sample $\overline{X}_{i}$ changing the indicator $c_{ij}$ to the opposite $\overline{c}_{ij}=1-c_{ij}$. Thus, $\hat{S}_{1}^{C}(t)$ and $\hat{S}^{C}_{2}(t)$ are the Kaplan–Meier estimates for additional samples $\overline{X}_{1}$ and $\overline{X}_{2}$.
The test statistic [32] can be computed as $S_{WKM}=U/\sqrt{\sigma^{2}}$ where
[…]
and $T_{C}=\max_{t\in R}\left(\min\{\hat{S_{1}}(t),\hat{S_{2}}(t),\hat{S_{1}}^{C}(t),\hat{S_{2}}^{C}(t)\}>0\right).$
The weight function $w(t)$ downweights the contributions of the $S_{2}(t)-S_{1}(t)$ in $U$ over later time periods if censoring is heavy. The statistic $U$ is asymptotically normal distributed with a variance $\sigma^{2}$:
[…]
The test statistic $S_{WKM}$ is asymptotically distributed as the standard normal distribution $N(0,1)$ with two-side critical area.
3.2.4 Bagdonavičius-Nikulin tests
Test based on the Simple Cross-Effect (SCE) model. The test statistic [23] is $S_{\text{BN-SCE}}=(U_{1},U_{2})^{T}\Sigma^{-1}(U_{1},U_{2})$ where
$U_{1}=\sum_{j=1}^{n_{1}}(1-c_{1j})\frac{K_{1}(t_{1j})}{Y(t_{1j})}Y_{2}(t_{1j})-\sum_{j=1}^{n_{2}}(1-c_{2j})\frac{K_{1}(t_{2j})}{Y(t_{2j})}Y_{1}(t_{2j}),$
$U_{2}=\sum_{j=1}^{n_{1}}(1-c_{1j})\frac{K_{2}(t_{1j})}{Y(t_{1j})}Y_{2}(t_{1j})-\sum_{j=1}^{n_{2}}(1-c_{2j})\frac{K_{2}(t_{2j})}{Y(t_{2j})}Y_{1}(t_{2j}).$
The elements of the matrix $\Sigma$ can be computed as
$\sigma_{ij}=\sum_{r=1}^{2}\sum_{s=1}^{n_{r}}(1-c_{rs})K_{i}(t_{rs})\frac{Y_{1}(t_{rs})Y_{2}(t_{rs})}{Y^{2}(t_{rs})}.$
The following functions are $K_{1}(t)$ and $K_{2}(t)$ may be used
$K_{1}(t)=exp(-\Lambda(t))/\sqrt{n},K_{2}(t)=(exp(-\Lambda(t))-1)/\sqrt{n},$
$Y(t)=Y_{1}(t)+Y_{2}(t),Y_{i}(t)=\sum_{j=1}^{n_{i}}Y_{ij}(t),Y_{ij}(t)=1_{t_{ij}\geq t},i=1,2,$
$\Lambda(t)=\sum_{i=1}^{2}\sum_{j:c_{ij}-0,t_{ij}\leq t}\frac{1}{Y(t_{ij})}.$
The test statistic $S_{\text{BN-SCE}}$ is asymptotically distributed as chi-square distribution with two degrees of freedom. The test has right-side critical area.
Test based on the Multiple Cross-Effect (MCE) model. The test statistic [24] is $S_{\text{BN-MCE}}=(U_{1},U_{2},U_{3})^{T}\Sigma^{-1}(U_{1},U_{2},U_{3}),$ where
$U_{1}=\sum_{j:c_{1j}=0}\frac{Y_{2}(t_{1j})}{Y(t_{1j})}-\sum_{j:c_{2j}=0}\frac{Y_{1}(t_{2j})}{Y(t_{2j})},U_{2}=-\sum_{j:c_{1j}=0}\frac{Y_{2}(t_{1j})}{Y(t_{1j})}\Lambda(t_{1j})-\sum_{j:c_{2j}=0}\frac{Y_{1}(t_{2j})}{Y(t_{2j})}\Lambda(t_{2j}),$
$U_{3}=-\sum_{j:c_{1j}=0}\frac{Y_{2}(t_{1j})}{Y(t_{1j})}\Lambda^{2}(t_{1j})-\sum_{j:c_{2j}=0}\frac{Y_{1}(t_{2j})}{Y(t_{2j})}\Lambda^{2}(t_{2j}).$
The elements of the matrix $\Sigma$ can be computed as
$\sigma_{11}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}\frac{Y_{1}(t_{ij})Y_{2}(t_{ij})}{Y^{2}(t_{ij})},\sigma_{12}=\sigma_{21}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}\frac{Y_{1}(t_{ij})Y_{2}(t_{ij})}{Y^{2}(t_{ij})}\Lambda(t_{ij}),$
$\sigma_{13}=\sigma_{31}=\sigma_{22}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}\frac{Y_{1}(t_{ij})Y_{2}(t_{ij})}{Y^{2}(t_{ij})}\Lambda^{2}(t_{ij}),$
$\sigma_{23}=\sigma_{32}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}\frac{Y_{1}(t_{ij})Y_{2}(t_{ij})}{Y^{2}(t_{ij})}\Lambda^{3}(t_{ij}),\sigma_{33}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}\frac{Y_{1}(t_{ij})Y_{2}(t_{ij})}{Y^{2}(t_{ij})}\Lambda^{4}(t_{ij}).$
The test statistic $S_{\text{BN-MCE}}$ is asymptotically distributed as chi-square distribution with three degrees of freedom. The test has right-side critical area.
Test based on the Generalized Proportional Hazards (GPH) model. The test statistic [25] is $S_{\text{BN-GPH}}=(U_{1},U_{2})^{T}\Sigma^{-1}(U_{1},U_{2})$ where
$U_{1}=\sum_{j:c_{1j}=0}\frac{Y_{2}(t_{1j})}{Y(t_{1j})}-\sum_{j:c_{2j}=0}\frac{Y_{1}(t_{2j})}{Y(t_{2j})},$
$U_{2}=-\sum_{j:c_{1j}=0}\frac{Y_{2}(t_{1j})}{Y(t_{1j})}\ln(1+\Lambda(t_{1j}))+\sum_{j:c_{2j}=0}\frac{Y_{1}(t_{2j})}{Y(t_{2j})}\ln(1+\Lambda(t_{2j})).$
The elements of the matrix $\Sigma$ can be computed as
$\sigma_{11}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}Y_{1}(t_{ij})Y_{2}(t_{ij})/Y^{2}(t_{ij}),$
$\sigma_{22}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}Y_{1}(t_{ij})Y_{2}(t_{ij})\ln^{2}(1+\Lambda(t_{ij}))/Y^{2}(t_{ij}),$
$\sigma_{12}=\sigma_{21}=\sum_{i=1}^{2}\sum_{j:c_{ij}=0}Y_{1}(t_{ij})Y_{2}(t_{ij})\ln(1+\Lambda(t_{ij}))/Y^{2}(t_{ij}).$
The test statistic $S_{\text{BN-GPH}}$ is asymptotically distributed as chi-square distribution with two degrees of freedom. The test has right-side critical area.
3.2.5 Two-Stage tests
Q-test. The test statistic [4] is $S_{Q}=S_{LG}$ if $Q<0$ else $S_{P}$ where $Q=\{\hat{S_{2}}(q_{0.6})-\hat{S_{1}}(q_{0.6})\}-\{\hat{S_{2}}(q_{0.2})-\hat{S_{1}}(q_{0.2})\}$, $q_{p}=\hat{S_{1}}^{-1}(p)$, and $\hat{S_{i}}(t)$ is the Kaplan-Meier estimator.
The test statistic $S_{Q}$ has a double-side critical area and the $H_{0}$ is rejected under a large value of the $|S_{Q}|$.
Maximum value test. The test statistic [26] is $S_{MAX}=\max(|S_{LG}|,|S_{G}|)$.
The test statistic $S_{MAX}$ has right-side critical area. The distribution of the test statistic $S_{MAX}$ under the $H_{0}$ is represented in [33].
MIN3 test. The test statistic [27] is $S_{MIN3}=\min(pv_{WKM},pv_{\text{BN-MCE}},pv_{\text{BN-GPH}})$ where $pv_{WKM}=2\cdot\min\{F_{N(0,1)}(S_{WKM}),1-F_{N(0,1)}(S_{WKM})\},$ $pv_{\text{BN-MCE}}=1-F_{\chi^{2}(3)}(S_{\text{BN-MCE}})$, $pv_{\text{BN-GPH}}=1-F_{\chi^{2}(2)}(S_{\text{BN-GPH}})$.
The test statistic $S_{MIN3}$ has left-side critical area. The distribution of the test statistic $S_{MIN3}$ under the $H_{0}$ is represented in [27].
3.3 Alternative Hypotheses
We study the performance of two-sample tests using alternative hypotheses where each alternative is a pair of competing hypotheses $S_{1}(t)$ and $S_{2}(t)$. Studies [5, 6] examine the impact of alternatives with differences at early, middle, and late time on the test power. Studies of Bagdonavičius and Nikulin [23, 24] use another way based on alternatives with different numbers of intersections of competing hypotheses. We combine these experiments together and provide experiments that have alternatives with 0, 1, and 2 intersections at early, middle, and late time.
Taking into account that the findings obtained for one alternative are not representative in the general case. Therefore, we form groups of similar alternatives that contain various closest competing hypotheses. The alternative groups are represented in Figure 1. Additionally, in order to reduce the impact of a random choice for alternative, we vary families of $F(t)$ and $F^{C}(t)$ for each alternative. The following families of distributions are used in the construction of alternative hypotheses:
  1. 1.
    $f_{Exp}(x;\mu,\lambda)=\lambda\exp(-\lambda(x-\mu)),x\geq\mu$;
  2. 2.
    $f_{We}(x;\mu,\lambda,\nu)=\nu(x-\mu)^{\nu-1}\exp(-((x-\mu)/\lambda)^{\nu})/\lambda^{\nu},x\geq\mu$;
  3. 3.
    $f_{G}(x;\mu,\lambda,\nu)=((x-\mu)/\lambda)^{\nu-1}\exp(-(x-\mu)/\lambda)\lambda\Gamma(\nu),x\geq\mu$;
  4. 4.
    $f_{LgN}(x;\mu,\lambda)=\exp(-(\ln{x}-\mu)^{2}/2\lambda^{2})/x\sqrt{2\pi\lambda^{2}},x\geq 0$.
Figure 1: Groups of alternative hypotheses
Note that we fix the size of inequality between competing hypotheses $S_{1}(t)$ and $S_{2}(t)$ by $L_{1}=0.1$ where $L_{1}$ is the $l1$-norm. It is necessary to ensure the closest experimental conditions between alternatives. The choice of this value is due to the following reasons:
  1. 1.
    It provides similar conditions between alternatives to ensure numerical experiments.
  2. 2.
    It provides the test power $1-\beta$ to be in the range $\alpha<1-\beta<1$ for most considered sample sizes where $\alpha$ and $\beta$ are errors of the I and II types.
[…]
Table 1: Groups of alternative hypotheses
Group
Description
$H_{i}$ $S_{1}(t)$ $S_{2}(t)$
I 0 intersections with difference in the early time $H_{01}$ $Exp(0,1)$ $Exp(0.1,1)$
$H_{02}$ $We(0,1.1,2.4)$ $LgN(0,0.370)$
$H_{03}$ $LgN(0.01,0.913)$ $Exp(0,0.742)$
II 0 intersections with difference in the middle time $H_{04}$ $G(0,1.060,1.160)$ $Exp(0,0.863)$
$H_{05}$ $Exp(0,1.3)$ $We(0,0.9,1.1)$
$H_{06}$ $Exp(0,1)$ $We(0.09,1.1,1.07)$
III 0 intersections with difference in the late time $H_{07}$ $Exp(0,1.3)$ $G(0,0.806,1.064)$
$H_{08}$ $We(0.5,1,1.2)$ $Exp(0.567,1)$
$H_{09}$ $We(0.118,1.1,1.735)$ $LgN(0.01,0.6)$
IV 1 intersection in the early time $H_{11}$ $Exp(0,1)$ $Exp(0.05,1.159)$
$H_{12}$ $G(0,1.273,1.475)$ $G(0.159,1.300,1.273)$
$H_{13}$ $We(0.02,1,1.1)$ $Exp(0,0.909)$
V 1 intersection in the middle time $H_{14}$ $We(0,0.980,0.905)$ $G(0,0.972,0.974)$
$H_{15}$ $Exp(0,1)$ $We(0.071,0.906,1.059)$
$H_{16}$ $G(0.01,1,1.15)$ $Exp(0,0.833)$
VI 1 intersection in the late time $H_{17}$ $We(0,0.968,1.214)$ $Exp(0,1.107)$
$H_{18}$ $G(0,1.1,1.040)$ $G(0,0.9,1.302)$
$H_{19}$ $We(0.5,1.1,1.1)$ $Exp(0.471,1)$
VII 2 intersections in the early and middle time $H_{21}$ $LgN(0,0.948)$ $We(0.173,1.325,0.911)$
$H_{22}$ $Exp(0.5,1.047)$ $LgN(0.141,0.596)$
$H_{23}$ $We(0.5,1,1.2)$ $Exp(0.530,1)$
VIII 2 intersections in the early and late time $H_{24}$ $LgN(0,0.916)$ $G(0.01,1.213,1.192)$
$H_{25}$ $LgN(0,0.817)$ $Exp(0.185,0.818)$
$H_{26}$ $We(0.01,1.697,1.846)$ $LgN(0.293,0.569)$
IX 2 intersections in the middle and late time $H_{27}$ $We(0,1.355,1.018)$ $LgN(0.000,0.867)$
$H_{28}$ $G(0,1.134,1.231)$ $LgN(0,0.876)$
$H_{29}$ $Exp(0,0.744)$ $LgN(0,0.866)$
4 Proposed ML-based Methods for Two Sample Problem
We approach the two-sample problem as a binary classification task (0 if the $H_{0}$ is true, 1 if the $H_{0}$ is rejected). Each two-sample test can be considered as an independent feature. Therefore, to solve our problem, we combine the properties of various statistical methods using a stacking architecture. For the binary classification task, we consider various machine learning methods:
  1. 1.
    Logistic Regression (denote as LogReg) implemented in sklearn;
  2. 2.
    Random Forest (denote as RF) implemented in sklearn;
  3. 3.
    Gradient Boosting Machine (denote as GBM) implemented in sklearn;
  4. 4.
    XGBoost;
  5. 5.
    CatBoost [34] proposed by Yandex;
  6. 6.
    LightAutoML [35] (denote as LAMA) proposed by Sber AI Lab.
Our method takes two right-censored samples $X_{1}$ and $X_{2}$ as input. For these samples, we compute the $p$-value for the two-sample tests described earlier, since small $p$-value rejects the null hypothesis $H_{0}$. We also include information about the sample sizes and the censoring rate in order to be more sensitive for the alternative. As output, our method returns a value in the interval $[0,1]$, which serves as the test statistic for a new two-sample test. For this method, we can estimate the test power, investigate the test statistic distribution under the null hypothesis $G(S|H_{0})$, and conduct any other experiments. We detail the process of the described method in Figure 2.
Figure 2: The implementation flow chart of the proposed methods
The test statistic of each proposed method has a right-side critical area (according to the problem statement based on the binary classification task) and the null hypothesis $H_{0}$ is rejected when the test statistic is large.
For the proposed methods, we limit the number of features at the feature engineering stage, since the goal is to add information to the methods about the properties of the two-sample tests on various sample sizes and censoring rates instead of overfitting on the selected alternatives.
In training ML models, we select the best models by maximizing the Accuracy metric. Since we have a balanced dataset (50% with the $H_{0}$ is true and 50% with the $H_{0}$ is rejected) of synthetic data, maximizing Accuracy allows us to minimize the probabilities of Type I and Type II errors. The synthetic data were simulated using the Monte Carlo replications. Also, we report other metrics for reference: the Precision, Recall, ROC AUC, Average Precision, Specificity and Proportion of predictions without FN (false negative): $(TP+FP+TN)/(TP+FP+FN+TN)$.
You also can solve the two-sample problem by the proposed methods which are available in GitHub.
5 Numerical Experiments
5.1 Dataset
To conduct numeric experiments for the proposed methods, we have simulated dataset including statistics of the two-sample tests described earlier. Each observation in the dataset contains the set of computed two-sample test statistics under following parameters:
  • 27 alternative hypotheses ($H_{01}-H_{09}$, $H_{11}-H_{19}$, and $H_{21}-H_{29}$);
  • 2 competing hypotheses $S_{1}(t)$ and $S_{2}(t)$ for each alternative:
    1. 1.
      if $X_{1}\sim\xi_{1}$ compares to $X_{2}\sim\xi_{1}$, then $H_{0}$ is true. It is indicated as ’H0’;
    2. 2.
      if $X_{1}\sim\xi_{1}$ compares to $X_{2}\sim\xi_{2}$, then $H_{0}$ is rejected. It is indicated as ’H1’;
  • 10 sample sizes for each competing hypothesis (20, 30, 50, 75, 100, 150, 200, 300, 500, and 1000 observations);
  • 6 censoring rates for each competing hypothesis (0%, 10%, 20%, 30%, 40%, 50%);
  • 37,650 Monte Carlo replications for each competing hypothesis.
For each observation in the dataset, the target variable indicates ’H0’ if the null hypothesis is true, or ’H1’ if the null hypothesis is rejected. In total, the dataset includes 27 x 2 x 10 x 6 x 37,650 = 121,986,000 observations. We split the dataset into two samples:
  1. 1.
    7,650 Monte Carlo replications with all combinations of parameters ($\sim 24.7$ million observations) to train the proposed methods (Train sample);
  2. 2.
    rest of 30,000 Monte Carlo replications (Simulation sample) for statistical investigations (test power, test statistic distributions and etc.).
This full dataset, split into samples, source code (C++) for simulation, and a brief summary are available in the Hugging Face repository (DOI: 10.57967/hf/2978).
5.2 Proposed Methods Training
In this numeric experiment, we train the proposed methods, then test them using several metrics and assess the feature importance. To reach this, we stratify the Train sample into: train subsample (55%) with $\sim 13.6$ million observations, validate subsample (30%) with $\sim 7.4$ million observations, and test subsample (15%) with $\sim 3.7$ million observations.
The ratio of class balance is equal to 50% in each sample. To find optimal hyperparameters, we apply the Optuna framework with 10 trials. The resulting metrics computed on the test sample are presented in Table 2.
Table 2: Training results of the proposed methods (with 95%-th confidence intervals)
Metric LogReg RF GBM
Accuracy 0.633 ± 0.0006 0.6459 ± 0.0007 0.6465 ± 0.0008
Precision 0.6305 ± 0.0007 0.7012 ± 0.0011 0.6955 ± 0.0010
Recall 0.6422 ± 0.0005 0.5086 ± 0.0008 0.5211 ± 0.0010
ROC AUC 0.686 ± 0.0009 0.7155 ± 0.0005 0.7153 ± 0.0006
Avg. Precision 0.6945 ± 0.0011 0.743 ± 0.0006 0.7423 ± 0.0007
Specificity 0.6237 ± 0.0010 0.7833 ± 0.0008 0.7719 ± 0.0007
Prop. w/o FN 0.8211 ± 0.0002 0.7543 ± 0.0004 0.7605 ± 0.0005
Metric XGBoost CatBoost LAMA
Accuracy 0.672 ± 0.0010 0.674 ± 0.0007 0.6771 ± 0.0008
Precision 0.7197 ± 0.0013 0.7181 ± 0.0012 0.7281 ± 0.0012
Recall 0.5635 ± 0.0008 0.5728 ± 0.0006 0.5653 ± 0.0007
ROC AUC 0.7478 ± 0.0009 0.7501 ± 0.0008 0.7539 ± 0.0008
Avg. Precision 0.7772 ± 0.0008 0.7789 ± 0.0008 0.7828 ± 0.0008
Specificity 0.7805 ± 0.0012 0.7751 ± 0.0012 0.7888 ± 0.0012
Prop. w/o FN 0.7818 ± 0.0004 0.7864 ± 0.0003 0.7827 ± 0.0003
The sequence of the most preferred methods based on the primary metric, Accuracy, is as follows: LAMA $>$ CatBoost $>$ XGBoost $>$ GBM or RF $>$ LogReg. This indicates that, on the test subsample, LAMA significantly reduces Type I and Type II errors compared to the other methods.
Next, let us analyze the Proportion without False Negatives (Type II error). In this scenario, the preferred methods rank as follows: LogReg $>$ CatBoost $>$ LAMA $>$ XGBoost $>$ GBM $>$ RF. This suggests that LogReg minimizes Type II errors more effectively than the other proposed methods on the test subsample.
Additionally, we discuss the ranking ability of the proposed methods as measured by Average Precision. This metric assesses the methods’ effectiveness in distinguishing differences in ordered values, particularly at the extremes of distributions. Given that our proposed methods focus on the right-side critical area, they demonstrate an ability to distinguish between competing hypotheses. The methods are ranked for this metric as follows: LAMA $>$ CatBoost $>$ XGBoost $>$ RF or GBM $>$ LogReg.
We also evaluate the importance of the features included in the final versions of each proposed method. For this, we use two methods: Permutation Importance (available for all methods) and Feature Importance (available for some methods, as an additional assessment). The results are shown in Tables 3 and  4.
Table 3: Permutation Importance (features sorted by average rank among all methods)
feature LogReg RF GBM XGBoost CatBoost LAMA
BN-SCE 21.24 0.95 1.64 18.92 26.81 12.03
BN-GPH 4.42 0.18 0.30 24.20 23.53 11.49
Tarone-Ware 1.01 0.15 0.05 27.66 26.75 12.72
MIN3 1.02 3.33 2.55 16.16 12.56 7.84
BN-MCE 15.59 2.15 3.59 6.58 7.64 4.41
WKM 8.22 0.06 0.02 21.37 16.55 10.70
Gehan 21.36 0.05 0.03 13.30 17.83 8.33
MAX 2.58 0.06 0.02 12.85 16.93 9.58
log-rank 0.08 0.13 0.15 12.81 13.20 7.00
Peto 1.59 0.04 0.00 17.40 7.21 8.70
Peto-Prentice 0.21 0.08 0.13 7.70 9.47 4.31
Prentice 1.29 0.10 0.07 2.38 3.76 3.12
$min(n_{1},n_{2})\geq 500$ 0.42 0.99 1.09 1.95 1.66 0.77
$Q$-test 0.53 0.03 0.04 5.83 4.28 4.91
$500<min(n_{1},n_{2})<$ 50 0.08 0.48 0.44 0.29 0.46 0.20
$500100\leq min(n_{1},n_{2})<$ 500 0.42 0.08 0.05 0.52 0.39 0.09
$10050\leq min(n_{1},n_{2})<$ 100 0.08 0.10 0.08 0.06 0.12 0.05
$\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}>35\%$ 0.68 0.00 0.00 0.27 0.35 0.24
$\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}=0\%$ 0.04 0.00 0.00 0.19 0.25 0.21
$15\%<\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}\leq 35\%$ 0.40 0.00 0.00 0.05 0.03 0.03
$0\%<\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}\leq 15\%$ 0.11 0.00 0.00 0.03 0.05 0.05
Table 4: Feature Importance (features sorted by average rank among all methods)
feature CatBoost XGBoost RF GBM
MIN3 20.08 40.16 24.37 17.81
$min(n_{1},n_{2})\geq 500$ 12.37 18.18 7.99 9.23
BN-MCE 7.33 4.43 20.44 29.32
BN-SCE 10.48 3.46 13.90 12.20
BN-GPH 8.88 2.07 10.01 15.50
$500<min(n_{1},n_{2})<$ 50 1.30 4.59 2.53 3.71
Tarone-Ware 8.33 1.57 1.46 0.87
log-rank 7.44 1.84 0.99 0.94
Gehan 2.68 1.04 3.17 1.27
$500100\leq min(n_{1},n_{2})<$ 500 1.86 9.16 1.14 0.85
$Q$-test 4.51 1.35 0.70 4.35
WKM 3.56 1.41 2.62 0.60
Peto 1.47 0.82 2.59 1.03
MAX 2.00 1.02 5.28 0.40
$10050\leq min(n_{1},n_{2})<$ 100 0.35 2.81 0.96 0.78
Prentice 2.39 0.84 1.02 0.52
Peto-Prentice 2.94 0.62 0.83 0.64
$\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}>35\%$ 1.05 1.55 0.00 0.00
$\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}=0\%$ 0.74 1.60 0.00 0.00
$0\%<\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}\leq 15\%$ 0.15 0.69 0.00 0.00
$15\%<\frac{n_{1}r_{1}+n_{2}r_{2}}{n_{1}+n_{2}}\leq 35\%$ 0.10 0.78 0.00 0.00
As a result, we present several key findings for the proposed methods, including: a) as expected in [27], one of most significant features are the two-sample Bagdonavičius-Nikulin and MIN3 tests; b) features based on the sample size are more significant than features based on the censoring rate; c) features of the censoring rate are at a level close to the insignificance.
5.3 Study of the Test Power
In this experiment, we determine the statistical power of the proposed methods. To do this, we compute estimates of the power of all the methods considered using the alternative hypotheses described earlier. As the result obtained, rather than reporting absolute power values, we compute the average rank among the set of methods considered, since the average rank more clearly reflects the preference of one method in comparison with rest methods. Detailed absolute values of the test power are represented in Supplementary materials.
For values of the average rank, we also compute the values of the Wald [36] and Savage [37] criteria, which are used for decision-making under risk and uncertainty of environment. In this case, the ”environment” is an alternative hypothesis, and the utility function is the rank of the method on the alternative hypothesis. The lower the rank of the method, the higher it is in the ordered list, which means that this method is better than others. The Wald criterion focuses on maximizing utility in worst-case scenarios, whereas the Savage criterion aims to minimize ’regret’ associated with the utility loss compared to the best possible outcome. A similar methodology has already applied previously in [38].
To reach this, we provide the investigation using the Simulation sample for each alternative group in the range $H_{I}$-$H_{IX}$ under censoring rates 0% and 0%-50%. We present results in Figure 3. The obtained results are ordered by the average rank values (AVG column).
Figure 3: Average rank (AVG) of tests including the Wald and Savage criteria on the groups $H_{\text{I}}-H_{\text{IX}}$
The results indicate that the proposed methods are at the top of the ordered lists (under censoring rate 0%-50%) with the exception of LogReg. This suggests that the proposed methods, on average, have a more preferable test power than the classic two-sample tests in terms of average rank. Comparing the values of the Wald criterion (assuming that the environment is in the worst state), we note that the proposed methods are also more preferable (under censoring rate 0%-50%). Nevertheless, the classic two-sample MIN3 test was among the TOP-4 preferred methods, since the construction of this test statistic is based on the same assumption. Comparing the values of the Savage criterion (’regret’ minimization), we note that the proposed methods also demonstrate superior performance (under censoring rate 0%-50%). Thus, we have shown that the proposed methods have high potential, since they have higher test power in the numerical experiments conducted.
5.4 Study of the Test Statistic Distribution under the $H_{0}$
Another investigation that is necessary for practical application is the study of the test statistic distribution. This ensures the accurate determination of the critical value in hypothesis testing. As is known, the actual (pre-limit) distribution of statistics can differ from the limit distribution and depend on many factors, such as sample size, censoring rate, data distribution, and many others. Therefore, it is important to evaluate its behavior in a numerical experiment in order to make correct statistical conclusions in practice.
The design of our numeric experiment is as follows:
  1. 1.
    For each set of parameters $(n_{1},n_{2},r_{1},r_{2},F,F^{C})$, we plot the CDF of the proposed method under the $H_{0}$ is true. We briefly denote such a distribution as $G(S|H_{0})$. Each $G(S|H_{0})$ contains 30,000 observations of the Simulation sample;
  2. 2.
    Construct a minorant $G_{MIN}(S|H_{0})$, a majorant $G_{MAX}(S|H_{0})$ among the set of the CDFs (obtained in item 1), and the middle line between them: $G_{AVG}(S|H_{0})=(G_{MIN}(S|H_{0})+G_{MAX}(S|H_{0}))/2$;
  3. 3.
    Present the average deviation of the middle line from the minorant/majorant in the right tail with probabilities $p$ in the range [0.9, 1.0). This is related to the fact that the proposed methods have right-critical area. Hence, it corresponds to area of the test size $0<\alpha\leq 0.1$).
This way allow us to estimate the average error in the $p$-value computation and assess a confidence interval for the $p$-value when solving a practical task. Table 5 presents computed average errors varing the proposed method, sample size, and censoring rate. Figure 4 shows the same results graphically.
Table 5: Average error of the $p$-value computation simulated for the test size $0\leq\alpha\leq 0.1,H_{01}-H_{29}$
$r_{i}$ $n_{i}$ LogReg RF GBM XGBoost CatBoost LAMA
0% 20 0.0053 0.0141 0.0136 0.0287 0.0231 0.0241
50 0.0079 0.0145 0.0084 0.0149 0.0166 0.0189
100 0.0084 0.0112 0.0078 0.0087 0.0088 0.0106
1000 0.0111 0.0100 0.0080 0.0069 0.0090 0.0083
10% 20 0.0077 0.0140 0.0114 0.0224 0.0284 0.0236
50 0.0072 0.0177 0.0099 0.0212 0.0218 0.0251
100 0.0091 0.0127 0.0086 0.0112 0.0114 0.0100
1000 0.0157 0.0122 0.0091 0.0170 0.0160 0.0133
20% 20 0.0105 0.0170 0.0159 0.0229 0.0224 0.0203
50 0.0057 0.0158 0.0096 0.0173 0.0209 0.0171
100 0.0075 0.0111 0.0092 0.0124 0.0148 0.0093
1000 0.0166 0.0107 0.0086 0.0177 0.0166 0.0099
30% 20 0.0105 0.0175 0.0164 0.0283 0.0319 0.0315
50 0.0055 0.0178 0.0112 0.0195 0.0206 0.0163
100 0.0090 0.0161 0.0150 0.0218 0.0245 0.0165
1000 0.0144 0.0122 0.0107 0.0176 0.0206 0.0105
40% 20 0.0122 0.0201 0.0189 0.0499 0.0521 0.0484
50 0.0070 0.0145 0.0114 0.0197 0.0223 0.0155
100 0.0074 0.0155 0.0138 0.0128 0.0186 0.0144
1000 0.0138 0.0114 0.0111 0.0205 0.0279 0.0128
50% 20 0.0160 0.0240 0.0231 0.0791 0.0882 0.0873
50 0.0064 0.0169 0.0123 0.0399 0.0421 0.0327
100 0.0081 0.0136 0.0117 0.0196 0.0225 0.0158
1000 0.0100 0.0115 0.0103 0.0339 0.0334 0.0216
The results indicate that cumulative distribution functions (CDFs) have the largest distance $|G_{MAX}(S|H_{0})-G_{MIN}(S|H_{0})|$ in the central area, and a significant decrease in distance occurs towards the tails of the distributions. Note that the proposed methods LogReg, RF and GBM have a smaller range between CDFs than XGBoost, CatBoost and LAMA. However, this is almost leveled out already when the sample size close to 100 observations. The deviations calculated in Table 5 suggest that the Monte Carlo simulation effectively addresses the determination of critical values in hypothesis testing. Thus, we have prepared percentage points which can be calculated for the proposed methods in order to use them in practice. Detailed values of percentage points are presented in Supplementary materials.
Figure 4: $G(S|H_{0})$ distributions for the proposed methods. White line is the middle line. Color stripe is the region between minorant and majorant.
6 Discussion
Thus, in this paper we proposed methods for two-sample testing under right-censored data based on various machine learning methods.
To train and test the proposed methods, we simulated a synthetic dataset (the dataset and source code are available in GitHub) containing over 121 million objects.
We trained the proposed methods on the training part of the dataset, computed metrics and determined which methods make the least number of type I and type II errors on the test part of the data (LAMA $>$ CatBoost $>$ XGBoost $>$ GBM or RF $>$ LogReg). The significance of features for the proposed methods based on Feature Importance and Permutation Importance was shown.
A power study was conducted for the proposed methods. On a set of close alternatives, we averaged the ranks of the methods, sorting them by test power. The lower the average rank, the more preferable the method is, on average. It was found that the proposed methods (especially LAMA and CatBoost) are more preferable methods for two-sample testing at a censoring rate of 0%-50%.
The behavior of the test statistic distribution of the proposed methods under the $H_{0}$ is true is also investigated. For this purpose, we constructed $G(S|H_{0})$ under various parameters $(n_{i},r_{i},F_{i},F^{C}_{i}),i=1,2$ and computed the average error between the middle line and the minorant/majorant of this set (in the region $0<\alpha\leq 0.1$). This shows what order of error can be expected when calculating the $p$-value. The obtained results show that Monte Carlo simulating the test statistic distribution to determine the critical value can be an effective way to solve this issue. A detailed table of percentage points was prepared.
The advantage of the proposed methods is that it is based on many classical and well-studied two-sample tests. This allows the method to be sensitive to the properties of these tests and aggregate their synergistic effect. Similar approaches based on the selection function have been used previously [4, 26, 27]. The choice of such a function was associated with taking into account the special statistical properties of the methods, which was a limitation for a wide study in this direction. ML methods allow one to be more sensitive to the properties of the underlying methods and at the same time play the role of the selection function. Note also that the obtained test power results confirm the results in [27] when studying the MIN3 test.
To apply these methods in practice, the full source code for the implementation of the proposed methods is published in the GitHub repository (including trained models and an example to run). If even this is not enough, the same repository contains the full source code for modeling the original dataset.
The proposed methods can find its application in solving problems in a variety of fields: medicine, biology, sociology, engineering problems and many others, where a comparison of two samples of right-censored observations (as well as complete observations) is required. However, for example, in genetics problems, the application of the proposed methods can be complicated by the lack of the limit distribution in an analytical form for calculating $p$-value at very low values of the type I error (e.g., $\alpha<0.001$). Nevertheless, this issue can also be resolved by increasing the volume of the Monte Carlo simulation.
However, one should not forget about the limitations of this work, which are as follows: (1) $n_{1}=n_{2}$ and $r_{1}=r_{2}$. It is necessary to check how the proposed methods behave under conditions of unequal sample sizes and/or censoring rates; (2) $n_{1},n_{2}\leq 1000$. Often, classical two-sample tests already have high power at such sample sizes. However, the behavior of the proposed methods for sample sizes greater than 1000 observations has not been investigated; (3) $r_{1},r_{2}\leq 50\%$. The extremely high level of censoring rate ($r>50\%$) requires additional investigation; (4) The set of alternative hypothesis has been various but still limited.
The study the limit test statistic distribution of the proposed methods can significantly increase the attractiveness and accessibility of these methods for solving practical problems. In our opinion, this can be an one of the prioritiest task for further possible research.
To summarize all of the above, we note that the use of the machine learning paradigm has already allowed us to qualitatively improve the solution of various applied problems. And this study also shows that the use of machine learning methods can significantly improve the accuracy of statistical inferences when two-sample problem testing.
To conduct the study, the following hardware is utilized: Intel® Core™ i7-12700H, 64GB RAM, 1TB ROM, and NVIDIA GeForce RTX 3060 Mobile.
7 Conclusion
In this study, we explored how machine learning methods can be applied to solve two-sample problems with right-censored observations. We formulated the solution to the problem as a stacking of classical two-sample tests (in fact, using machine learning methods as a selection function). To train and test this method, we prepared a synthetic dataset using Monte Carlo simulations.
Despite the lack of the limit distribution for calculating the $p$-value (this issue can be resolved using the Monte Carlo simulation), the results indicate that the proposed methods enhance the accuracy of statistical inferences in two-sample testing, whether under right-censored data or not. Testing this hypothesis is an important task in many areas, such as medicine, sociology, engineering, and many others.
To be able to reproduce the results and use the proposed methods, we published the original synthetic dataset, including the source code for modeling (Hugging Face), and the trained models, including the source code for running (GitHub).
Supplemental Materials
The supplementary materials contain: (1) Table with absolute values of simulated test power; (2) Table with percentage points of the proposed methods including an averaged error of the $p$-value.
Data Availability Statement
The dataset simulated during and/or analyzed during the current study is available in the Hugging Face repository (pfilonenko/ML_for_TwoSampleTesting/), DOI: 10.57967/hf/2978.
The source code for the dataset simulation is available in the GitHub repository (pfilonenko/ML_for_TwoSampleTesting/dataset/simulation/).
The source code for the proposed methods and trained models are available in the GitHub repository (pfilonenko/ML_for_TwoSampleTesting).
Disclosure Statement
The authors have no conflicts of interest to declare.
Funding
This research received no specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
ORCID